Sisters
Fran is 8 years older than Maria. Four years ago Fran was five times as old as Maria. How old are the girls?

Answer by macston(5194)

.
F=Fran's age; M=Maria's age
F=M+8 Use this to substitute for F.
.
F-4=5(M-4) Substitute for F
(M+8)-4=5(M-4)
M+4=5M-20
24=4M
6=M ANSWER 1: Maria is 6 years old.
.
F=M+8=6+8=14 ANSWER 2: Fran is 14 years old.
.
CHECK:
F-4=5(M-4)
14-4=5(6-4)
10=5(2)
10=10